# Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone

### v3

@@ Abstract
We introduce phi-3-mini, a 3.8 billion parameter language model trained on 3.3 trillion tokens, whose overall performance, as measured by both academic benchmarks and internal testing, rivals that of models such as Mixtral 8x7B and GPT-3.5 (e.g., phi-3-mini achieves 69% on MMLU and 8.38 on MT-bench), despite being small enough to be deployed on a phone. The innovation lies entirely in our dataset for training, a scaled-up version of the one used for phi-2, composed of heavily filtered publicly available web data and synthetic data. The model is also further aligned for robustness, safety, and chat format. We also provide some initial parameter-scaling results with a 7B and 14B models trained for 4.8T tokens, called phi-3-small and phi-3-medium, both significantly more capable than phi-3-mini (e.g., respectively 75% and 78% on MMLU, and 8.7 and 8.9 on MT-bench). Moreover, we also introduce phi-3-vision, a 4.2 billion parameter model based on phi-3-mini with strong reasoning capabilities for image and text prompts.
@@ 2 Technical Specifications
The phi-3-small model (7B parameters) leverages the tiktoken tokenizer (for better multilingual tokenization) with a vocabulary size of 100352222We remove unused tokens from the vocabulary. and has default context length $8192$. It follows the standard decoder architecture of a 7B model class, having $32$ heads, $32$ layers and a hidden size of $4096$. We switched from GELU activation to GEGLU and used Maximal Update Parametrization (muP) [yang2022tensor] to tune hyperparameters on a small proxy model and transfer them to the target 7B model. Those helped ensure better performance and training stability. Also, the model leverages a grouped-query attention, with $4$ queries sharing $1$ key. To optimize the training and inference speed, we design a novel blocksparse attention module. For each attention head, the blocksparse attention enforces different sparsity patterns over KV cache. This ensures that all tokens are attended to on different heads for the given choice of sparsity. As illustrated in Figure 1, the context is then efficiently divided and conquered among attention heads, with significant KV cache reduction. To achieve actual deployment speed-up from the blocksparse design, we implemented highly efficient, yet flexible kernels for both training and inference. For training, we build a triton kernel based on Flash Attention [DFE+22]. For inference, we implemented a kernel for the prefilling phase and extended the paged attention kernel in vLLM for the decoding phase [KLZ+23]. Lastly, in phi-3-small architecture, we alternate dense attention layers and blocksparse attention layers to optimize KV cache savings while maintaining long context retrieval performance. An additional 10% multilingual data was also used for this model.
@@
Post-training of phi-3-mini went through two stages, including supervised finetuning (SFT) and direct preference optimization (DPO). SFT leverages highly curated high-quality data across diverse domains, e.g., math, coding, reasoning, conversation, model identity, and safety. The SFT data mix starts with using English-only examples. DPO data covers chat format data, reasoning, and responsible AI (RAI) efforts. We use DPO to steer the model away from unwanted behavior, by using those outputs as “rejected” responses. Besides improvement in math, coding, reasoning, robustness, and safety, post-training transforms a language model to an AI assistant that users can efficiently and safely interact with.
As part of the post-training process, we developed a long context version of phi-3-mini with context length limit enlarged to 128K instead of 4K. Across the board, the 128K model quality is on par with the 4K length version, while being able to handle long context tasks. Long context extension has been done in two stages, including long context mid-training and long-short mixed post-training with both SFT and DPO.
@@ 3 Academic benchmarks
On the next page we report the results for phi-3-mini on standard open-source benchmarks measuring the model’s reasoning ability (both common sense reasoning and logical reasoning). We compare to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Mixtral-8x7b [JSR+24], Gemma 7B [TMH+24], Llama-3-instruct-8b [AI], and GPT-3.5. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable. These numbers might differ from other published numbers due to slightly different choices in the evaluation. As is now standard, we use few-shot prompts to evaluate the models, at temperature 00. The prompts and number of shots are part of a Microsoft internal tool to evaluate language models, and in particular we did no optimization to the pipeline for the phi-3 models. [^4] For example, we found that using ## before the Question can lead to a noticeable improvement to phi-3-mini’s results across many benchmarks, but we did not do such changes in the prompts. The number of $k$–shot examples is listed per-benchmark. An example of a 2-shot prompt is described in Appendix A.
@@
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21] 68.8 75.7 78.0 56.3 61.7 63.6 66.5 70.5 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 77.0 82.4 53.6 58.5 49.8 71.1 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 58.1 55.8 42.5 47.1 48.7 57.3 55.2 58.1 GSM-8K (8-Shot; CoT) [CKB+21] 82.5 89.6 91.0 61.1 46.4 59.8 77.4 64.7 78.1 MedQA (2-Shot) [JPO+20] 53.8 65.4 69.9 40.9 50.0 49.6 60.5 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.1 50.2 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 58.1 73.9 45.2 75.2 72.3 67.7 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.6 75.9 78.6 78.3 82.8 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.0 97.7 88.5 90.6 91.4 93.4 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 86.9 87.9 60.2 77.7 78.1 75.7 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.2 80.2 68.3 74.6 65.5 73.9 75.9 68.3 BigBench-Hard (3-Shot; CoT) [SRR+22, SSS+22] 71.7 79.1 81.4 59.4 57.3 59.6 51.5 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 81.5 81.5 54.7 54.2 55.6 65.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.0 87.4 73.6 79.8 78.6 82.6 85.8 86.0 BoolQ (2-Shot) [CLC+19] 77.2 84.8 86.5 72.2 66.0 80.9 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.0 82.8 69.3 72.6 76.2 79.0 78.1 79.6 TruthfulQA (10-Shot; MC2) [LHE22] 65.0 70.2 75.1 53.0 52.1 63.2 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 58.5 61.0 62.2 59.0 28.0 34.1 60.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.7 75.2 60.6 50.8 51.5 67.7 60.2 77.8 Average 71.2 75.7 78.5 61.2 61.7 69.4 69.8 74.3 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
@@ 4 Safety
The safety alignment of phi-3-small and phi-3-medium was conducted by undergoing the same red-teaming process, utilizing identical datasets, and incorporating a slightly larger number of samples. Table 1 shows the results of in-house RAI benchmarks [MHJ+23] for phi-3 models compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24], and Llama-3-instruct-8b [AI]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Ungroundedness 0.603 0.299 0.213 1.481 0.935 0.679 0.328 Third Party Harm (DR-1) 0.240 0.253 0.251 0.240 0.562 0.383 0.373 Harmful Content Continuation (DR-3) 0.007 0.003 0.010 0.029 0.026 0.013 0.013 Harmful Content Summarization (DR-3) 0.100 0.110 0.112 0.144 0.223 0.103 0.082 Jailbreak (DR-1) 0.123 0.107 0.111 0.150 0.156 0.114 0.130
@@ 5 Weakness
6 Phi-3-Vision
Figure 6: The demo case shows Phi-3-Vision’s capability in natural image understanding and reasoning.

### v4

@@ Abstract
We introduce phi-3-mini, a 3.8 billion parameter language model trained on 3.3 trillion tokens, whose overall performance, as measured by both academic benchmarks and internal testing, rivals that of models such as Mixtral 8x7B and GPT-3.5 (e.g., phi-3-mini achieves 69% on MMLU and 8.38 on MT-bench), despite being small enough to be deployed on a phone. Our training dataset is a scaled-up version of the one used for phi-2, composed of heavily filtered publicly available web data and synthetic data. The model is also further aligned for robustness, safety, and chat format. We also provide parameter-scaling results with a 7B, 14B models trained for 4.8T tokens, called phi-3-small, phi-3-medium, both significantly more capable than phi-3-mini (e.g., respectively 75%, 78% on MMLU, and 8.7, 8.9 on MT-bench). To enhance multilingual, multimodal, and long-context capabilities, we introduce three models in the phi-3.5 series: phi-3.5-mini, phi-3.5-MoE, and phi-3.5-Vision. The phi-3.5-MoE, a 16 x 3.8B MoE model with 6.6 billion active parameters, achieves superior performance in language reasoning, math, and code tasks compared to other open-source models of similar scale, such as Llama 3.1 and the Mixtral series, and on par with Gemini-1.5-Flash and GPT-4o-mini. Meanwhile, phi-3.5-Vision, a 4.2 billion parameter model derived from phi-3.5-mini, excels in reasoning tasks and is adept at handling both single-image and text prompts, as well as multi-image and text prompts.
@@ 2 Technical Specifications
The phi-3-small model (7B parameters) leverages the tiktoken tokenizer (for better multilingual tokenization) with a vocabulary size of 100352222We remove unused tokens from the vocabulary. and has default context length $8192$. It follows the standard decoder architecture of a 7B model class, having $32$ heads, $32$ layers and a hidden size of $4096$. We switched to GEGLU activation and used Maximal Update Parametrization (muP) [YHB+22] to tune hyperparameters on a small proxy model and transfer them to the target 7B model. Those helped ensure better performance and training stability. Also, the model leverages a grouped-query attention, with $4$ queries sharing $1$ key. To optimize the training and inference speed, we design a novel blocksparse attention module. For each attention head, the blocksparse attention enforces different sparsity patterns over KV cache. This ensures that all tokens are attended to on different heads for the given choice of sparsity. As illustrated in Figure 1, the context is then efficiently divided and conquered among attention heads, with significant KV cache reduction. To achieve actual deployment speed-up from the blocksparse design, we implemented highly efficient, yet flexible kernels for both training and inference. For training, we build a triton kernel based on Flash Attention [DFE+22]. For inference, we implemented a kernel for the prefilling phase and extended the paged attention kernel in vLLM for the decoding phase [KLZ+23]. Lastly, in phi-3-small architecture, we alternate dense attention layers and blocksparse attention layers to optimize KV cache savings while maintaining long context retrieval performance. An additional 10% multilingual data was also used for this model.
@@
The phi-3.5-MoE adopts an Mixture-of-Experts (MoE) architecture to selectively activate parts of modules on specific inputs to improve the model efficiency. It incorporates MoE layer as its feedforward models, employing the top2 routing among 16 expert networks. Particularly, each expert network is a separate GLU network and the routing module will selectively activate 2 expert networks out of the 16 expert networks for each token, leaving 16×3.8B model to have 6.6B activated parameters with 42B total parameters. Additionally, we utilize the SparseMixer approach [LGC23, LDL+23] for training the sparse router in the MoE model. For comparison with other Phi series models, phi-3.5-MoE uses the same tokenizer as phi-3-medium and phi-3-mini with vocabulary size of 32064.
@@
Post-training of phi-3 went through two stages, including supervised finetuning (SFT) and direct preference optimization (DPO). SFT leverages highly curated high-quality data across diverse domains, e.g., math, coding, reasoning, conversation, model identity, and safety. The SFT data mix starts with using English-only examples. DPO data covers chat format data, reasoning, and responsible AI (RAI) efforts. We use DPO to steer the model away from unwanted behavior, by using those outputs as “rejected” responses. Besides improvement in math, coding, reasoning, robustness, and safety, post-training transforms a language model to an AI assistant that users can efficiently and safely interact with.
@@ 3 Academic benchmarks
On the next page we report the results for phi-3 on standard open-source benchmarks measuring the model’s reasoning ability (both common sense reasoning and logical reasoning). We compare to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Mixtral-8x7b [JSR+24], Gemma 7B [TMH+24], Llama-3-instruct-8b [AI23], and GPT-3.5. All the reported numbers are produced with the exact same pipeline to ensure that the numbers are comparable. These numbers might differ from other published numbers due to slightly different choices in the evaluation. As is now standard, we use few-shot prompts to evaluate the models, at temperature 00. The prompts and number of shots are part of a Microsoft internal tool to evaluate language models, and in particular we did no optimization to the pipeline for the phi-3 models. [^4] For example, we found that using ## before the Question can lead to a noticeable improvement to phi-3-mini’s results across many benchmarks, but we did not do such changes in the prompts. The number of $k$–shot examples is listed per-benchmark. An example of a 2-shot prompt is described in Appendix A.
@@
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Mixtral 8x7b GPT-3.5 version 1106 MMLU (5-Shot) [HBK+21a] 68.8 75.7 78.0 56.3 61.7 63.6 66.5 70.5 71.4 HellaSwag (5-Shot) [ZHB+19] 76.7 77.0 82.4 53.6 58.5 49.8 71.1 70.4 78.8 ANLI (7-Shot) [NWD+20] 52.8 58.1 55.8 42.5 47.1 48.7 57.3 55.2 58.1 GSM-8K (8-Shot; CoT) [CKB+21] 82.5 89.6 91.0 61.1 46.4 59.8 77.4 64.7 78.1 MATH (0-Shot; CoT) [HBK+21b] 41.3 34.6 53.1 15.0 13.6 28.2 11.1 45.3 MedQA (2-Shot) [JPO+20] 53.8 65.4 69.9 40.9 50.0 49.6 60.5 62.2 63.4 AGIEval (0-Shot) [ZCG+23] 37.5 45.1 50.2 29.8 35.1 42.1 42.0 45.2 48.4 TriviaQA (5-Shot) [JCWZ17] 64.0 58.1 73.9 45.2 75.2 72.3 67.7 82.2 85.8 Arc-C (10-Shot) [CCE+18] 84.9 90.7 91.6 75.9 78.6 78.3 82.8 87.3 87.4 Arc-E (10-Shot) [CCE+18] 94.6 97.0 97.7 88.5 90.6 91.4 93.4 95.6 96.3 PIQA (5-Shot) [BZGC19] 84.2 86.9 87.9 60.2 77.7 78.1 75.7 86.0 86.6 SociQA (5-Shot) [BZGC19] 76.6 79.2 80.2 68.3 74.6 65.5 73.9 75.9 68.3 BigBench-Hard (3-Shot; CoT) [SRR+22, SSS+22] 71.7 79.1 81.4 59.4 57.3 59.6 51.5 69.7 68.32 WinoGrande (5-Shot) [SLBBC19] 70.8 81.5 81.5 54.7 54.2 55.6 65.0 62.0 68.8 OpenBookQA (10-Shot) [MCKS18] 83.2 88.0 87.4 73.6 79.8 78.6 82.6 85.8 86.0 BoolQ (2-Shot) [CLC+19] 77.2 84.8 86.5 72.2 66.0 80.9 77.6 79.1 CommonSenseQA (10-Shot) [THLB19] 80.2 80.0 82.8 69.3 72.6 76.2 79.0 78.1 79.6 TruthfulQA (10-Shot; MC2) [LHE22] 65.0 70.2 75.1 53.0 52.1 63.2 60.1 85.8 HumanEval (0-Shot) [CTJ+21] 58.5 61.0 62.2 59.0 28.0 34.1 60.4 37.8 62.2 MBPP (3-Shot) [AON+21] 70.0 71.7 75.2 60.6 50.8 51.5 67.7 60.2 77.8 Average 69.7 73.6 76.7 58.9 59.3 67.3 66.8 72.8 GPQA (2-Shot; CoT) [RHS+23] 32.8 34.3 29.0 MT Bench (2 round ave.) [ZCS+23] 8.38 8.70 8.91 8.35
4 Multilingual and Long Context
To enhance the Phi-3 models with multilingual and long-context capabilities, we developed the versions phi-3.5-mini and phi-3.5-MoE, which incorporate more multilingual and long-text data during mid-training. Specifically, we employed the long-rope method [DZZ+24a] and a mixed context window approach to expand the context length limit from 4K to 128K without compromising performance on 4K-context tasks.
Figure  4 compares the performance of phi-3-mini, phi-3.5-mini, and phi-3.5-MoE on MMLU multilingual tasks. phi-3.5-mini demonstrates significant improvement over phi-3-mini in languages such as Arabic, Chinese, Russian, Ukrainian, and Vietnamese, with average MMLU-multilingual scores of $55.4$ and $47.3$, respectively. Due to its larger model capacity, phi-3.5-MoE achieves a significantly higher average score of $69.9$, outperforming phi-3.5-mini.
Figure 4: Comparison of phi-3-mini, phi-3.5-mini and phi-3.5-MoE on MMLU-Multilingual tasks
We evaluate the phi-3.5-mini and phi-3.5-MoE models on two long-context understanding tasks: RULER [HSK+24] and RepoQA [LTD+24]. As shown in Tables 1 and 2, both phi-3.5-MoE and phi-3.5-mini outperform other open-source models with larger sizes, such as Llama-3.1-8B, Mixtral-8x7B, and Mixtral-8x22B, on the RepoQA task, and achieve comparable performance to Llama-3.1-8B on the RULER task. However, we observe a significant performance drop when testing the 128K context window on the RULER task. We suspect this is due to the lack of high-quality long-context data in mid-training, an issue we plan to address in the next version of the model release.
In the table 3, we present a detailed evaluation of the phi-3.5-mini and phi-3.5-MoE models compared with recent SoTA pretrained language models, such as GPT-4o-mini, Gemini-1.5 Flash, and open-source models like Llama-3.1-8B and the Mistral models. The results show that phi-3.5-mini achieves performance comparable to much larger models like Mistral-Nemo-12B and Llama-3.1-8B, while phi-3.5-MoE significantly outperforms other open-source models, offers performance comparable to Gemini-1.5 Flash, and achieves above 90% of the average performance of GPT-4o-mini across various language benchmarks.
Model Ctx Size Python C++ Rust Java TypeScript Average gpt-4O-2024-05-13 128k 95 80 85 96 97 90.6 gemini-1.5-flash-latest 1000k 93 79 87 94 97 90 Phi-3.5-MoE 128k 89 74 81 88 95 85 Phi-3.5-Mini 128k 86 67 73 77 82 77 Llama-3.1-8B-Instruct 128k 80 65 73 76 63 71 Mixtral-8x7B-Instruct-v0.1 32k 66 65 64 71 74 68 Mixtral-8x22B-Instruct-v0.1 64k 60 67 74 83 55 67.8
Table 1: Comparison results on RepoQA benchmark.
Model Ctx Size 4k 8k 16k 32k 64k 128k Average Llama-3.1-8B-Instruct 128k 95.5 93.8 91.6 87.4 84.7 77.0 88.3 Phi-3.5-MoE 128k 94.8 93.0 93.2 91.6 85.7 64.2 87.1 Phi-3.5-Mini 128k 94.3 91.1 90.7 87.1 78.0 63.6 84.1 Mixtral-8x22B-Instruct-v0.1 64k 95.6 94.9 93.4 90.9 84.7 31.7 81.9 Mixtral-8x7B-Instruct-v0.1 32k 94.9 92.1 92.5 85.9 72.4 44.5 80.4
Table 2: Comparison results on RULER benchmark.
Category Benchmark Phi-3.5-mini 3.8B Phi-3.5-MoE 16x3.8B Mistral 7B Mistral-Nemo 12B Llama-3.1-In 8B Gemma-2 9B Gemini-1.5 Flash GPT-4o-mini Popular Arena Hard 37 37.9 18.1 39.4 25.7 42 55.2 75 BigBench Hard CoT (0-shot) 69 79.1 33.4 60.2 63.4 63.5 66.7 80.4 MMLU MMLU (5-shot) 69 78.9 60.3 67.2 68.1 71.3 78.7 77.2 MMLU-Pro (0-shot, CoT) 47.5 54.3 18 40.7 44 50.1 57.2 62.8 Reasoning ARC Challenge (10-shot) 84.6 91.0 77.9 84.8 83.1 89.8 92.8 93.5 BoolQ (2-shot) 78 84.6 80.5 82.5 82.8 85.7 85.8 88.7 GPQA (0-shot, CoT) 27.2 36.8 15.6 28.6 26.3 29.2 37.5 41.1 HellaSwag (5-shot) 69.4 83.8 71.6 76.7 73.5 80.9 67.5 87.1 OpenBookQA (10-shot) 79.2 89.6 78 84.4 84.8 89.6 89 90 PIQA (5-shot) 81 88.6 73.4 83.5 81.2 83.7 87.5 88.7 Social IQA (5-shot) 74.7 78.0 73 75.3 71.8 74.7 77.8 82.9 TruthfulQA (10-shot,MC2) 64 77.5 64.7 68.1 69.2 76.6 76.6 78.2 WinoGrande (5-shot) 68.5 81.3 58.1 70.4 64.7 74 74.7 76.9 Multilingual Ml MMLU (5-shot) 55.4 69.9 47.4 58.9 56.2 63.8 77.2 72.9 MGSM (0-shot CoT) 47.9 58.7 31.8 63.3 56.7 76.4 75.8 81.7 Math GSM8K (8-shot, CoT) 86.2 88.7 54.4 84.2 82.4 84.9 82.4 91.3 MATH (0-shot, CoT) 48.5 59.5 19 31.2 47.6 50.9 38 70.2 Long context Qasper 41.9 40.0 31.4 30.7 37.2 13.9 43.5 39.8 SQuALITY 24.3 24.1 25.9 25.8 26.2 0 23.5 23.8 Code HumanEval (0-shot) 61.5 70.7 35.4 63.4 66.5 61 74.4 86.6 MBPP (3-shot) 68.6 80.8 50.4 68.1 69.4 69.3 77.5 84.1 Average 61.1 69.2 48.5 61.3 61.0 63.3 68.5 74.9
Table 3: Model quality on representative benchmarks
@@ 5 Safety
The safety alignment of phi-3-small, phi-3-medium and phi-3.5-MoE was conducted by undergoing the same red-teaming process, utilizing identical datasets, and incorporating a slightly larger number of samples. Table 4 shows the results of in-house RAI benchmarks [MHJ+23] for phi-3 models compared to phi-2 [JBA+23], Mistral-7b-v0.1 [JSM+23], Gemma 7b [TMH+24], and Llama-3-instruct-8b [AI23]. This benchmark utilized GPT-4 to simulate multi-turn conversations in five different categories and to evaluate the model responses. Ungroundedness between 0 (fully grounded) and 4 (not grounded) measures if the information in a response is based on a given prompt. In other categories, responses were evaluated in terms of the severity of harmfulness from 0 (no harm) to 7 (extreme harm) and the defect rates (DR-$x$) were computed as the percentage of samples with the severity score being greater than or equal to $x$.
Phi-3-mini 3.8b Phi-3-small 7b Phi-3-medium 14b Phi-3.5-MoE 16x3.8b Phi-2 2.7b Mistral 7b Gemma 7b Llama-3-In 8b Ungroundedness 0.603 0.299 0.213 0.228 1.481 0.935 0.679 0.328 Third Party Harm (DR-1) 0.240 0.253 0.251 0.105 0.240 0.562 0.383 0.373 Harmful Content Continuation (DR-3) 0.007 0.003 0.010 0.005 0.029 0.026 0.013 0.013 Harmful Content Summarization (DR-3) 0.100 0.110 0.112 0.12 0.144 0.223 0.103 0.082 Jailbreak (DR-1) 0.123 0.107 0.111 0.106 0.150 0.156 0.114 0.130
@@ 6 Weakness
7 Phi-3.5-Vision
Figure 7: The demo case shows Phi-3.5-Vision’s capability in natural image understanding and reasoning.